Barefoot Running Company's work in process inventory on June 1 has a balance of $22,900 representing Job No. 265. During June, $50,700 of direct materials were requisitioned for Job No. 265 and $35,500 of direct labor cost was incurred on Job No. 265. Manufacturing overhead is allocated at 130% of direct labor cost. Actual manufacturing overhead costs incurred in June amounted to $41,300. No new jobs were started during June. Job No. 265 is completed on June 28.

Is manufacturing overhead overallocated or underallocated for the month of June? By how much?

A.

$4,850 overallocated

B.

$4,850 underallocated

C.

$11,530 underallocated

D.

$11,530 overallocated
Answer #1

Overhead allocated on job 265 = 35500*130% = $46150

Actual overhead in June = $41300

Over allocated overhead = 46150-41300 = $4850

So answer is a) $4850 Over allocated